Please show all work X, be a random sample from the distribution with the probability density function Let A0 and let X, X2, f(x; A) 24xe, x>0. a.
Answer #1

2) f(0) -xx2 7 427812? xxo 7+k x2 Elak) Let x² = t. 27adx = dt. -t at Elxk) = 10 x 6tk 2x3 to 6 JoPage No. nos *in de We know that you e-x xl de We know O2 - that te 2 al. (e) = I 14+R2 b) E(x) = put k=1 in (a) we have 6/02Page No. 32/ m. 2 35 32 / 35 (32) 2-53 x 22 7 x 2² be = Ñ = 0.5874 L= L = taxi f(xi, e) n dun (3) - txi e- Am Exit & const. 8